If excess of water is added to 100g oleum sample labelled as 112% H2SO4 and is reacted with 5.3g Na2SO3,then find volume of CO2 evolved at 1 atm pressure and 300 K temperature after the reaction is complete.

Grade:12

1 Answers

Sunil Kumar FP
9 years ago
SO3 + H2O---H2SO4
18 g cof h20 combine with 80 g of so3
12 g combine with 160/3=53.3g
H2SO4 =100-53.3g=46.6g=.47mole
now
mole of Na2SO3=5.3/126=.042g
Na2SO3 +H2SO4----Na2So4 +SO2 +H20
mole of SO2=.042
VOL=nRT/P=.042*.083*300=1lt
thanks and regards
sunil kr
askIITian faculty